OpenStudy (anonymous):

Does anyone a SIMPLE example of an arbitrary constant? I've been looking for a few, but I keep getting confused, so can someone please give me a simple one?

OpenStudy (amistre64):

an arbitrary constant .... is just some constant that doesnt need to be determined at this moment

OpenStudy (anonymous):

@amistre64 Thanks, but do you have any example? Or could you explain this one: f(x)+k when f(x)=x+9×3, f(x)=x×6+9-3, and f(x)=x-32 for example, k is an arbitrary constant, because you add k to whatever function you have.

OpenStudy (amistre64):

either the sites not decodinng the encoding, or its my laptop ... i get black diamonds with a ? in them

OpenStudy (amistre64):

is this in calculus class by chance?

OpenStudy (anonymous):

oh that is a multiplication sign

OpenStudy (anonymous):

and no, this is a question that was supposed to be on my report for number systems

OpenStudy (amistre64):

hmm, im thinking of how it is defined for a calculus course. so the context might change its meaning a little

OpenStudy (amistre64):

ideally, an arbitrary constant allows us to define a 'family' of similar functions that differ only by a constant

OpenStudy (anonymous):

right, is there any simple example of this?

OpenStudy (amistre64):

outside of calculus, im not sure

OpenStudy (anonymous):

that's okay, thanks though:)

OpenStudy (amistre64):

an example, how simple it is i dunno what are ALL the solutions to the equation y' = 2x? well, y=x^2, or y=x^2+1 or y=x^2 + pi or y=x^2 + (some constant) ALL the solutions will be of the form: y = x^2 + (some constant) the constant is arbitrary as far as a general solution goes

OpenStudy (amistre64):

other than that, idk :)

OpenStudy (anonymous):

@amistre64 Thanks :)